If you invest $125 in a stock, borrowing 80 percent of the $125 at 20 percent interest, and the stock price rises by 40 percent, what is the return on your investment?

________ percent
Answer #1

Investment =$125.

Borrowed =0.8*125 = 100.

Interest paid = 0.2*100 =20.

Total amount to be repaid = 100+20 =120.

Return on investment = 125*0.4 = 50.

Total amount received =125+50 =175.

Return = 175-125-20 = 30.

Return on investment=30/125 = 24%.